# appendixReferences

### v3

@@ 4.2 Variance Estimation for General Case
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{X\in R_{i},Z_{i}=Y_{ki}\}}{P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})}$ (37)
so that $\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}v_{i}$. It follows that the variance of $\hat{\Phi}_{k}^{\textup{Haj}}$ can be written as a sum of covariances
${\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}{\operatorname{Cov}}(v_{i},v_{j})$
Unlike the special case where $T_{i}$ is constant for all units, generally we cannot know or compute ${\operatorname{Cov}}(v_{i},v_{j})$ without knowledge of the outcome models $f_{i}$ and $f_{j}$, even if $\phi_{i}^{*}\phi_{j}^{*}=1$. As a result, the covariance terms must be estimated. We thus propose variance estimator $V_{k}$ given by
$V_{k}=\sum_{i:\phi_{i}^{*}=1}\sum_{j:\phi_{j}^{*}=1}\Bigg{\{}\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}\left[v_{i}|T_{i}\right]\cdot\mathbb{E}\left[v_{j}|T_{j}\right]\cdot\\ \frac{P(T_{i})\cdot P(T_{j})}{P(T_{i},T_{j})}\Bigg{\}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}$ (38)
where $\bar{Z}_{i}$ equals $Z_{i}$ if $X\in R_{i}$ and equals -1 otherwise,
@@
$V_{k}=\sum_{i:\phi^{*}=1}\sum_{j:\phi^{*}=1}{\operatorname{Cov}}(v_{i},v_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})},$
@@
Theorem 3 gives computable expressions for the terms appearing in (38). Using these expressions, we may evaluate $V_{k}(\phi)$, the value of $V_{k}$ under the hypothesis that $\phi^{*}=\phi$, without knowledge of the true outcome models $\{f_{i}\}$.
Theorem 3.
If $\phi_{i}^{*}=1$ and $\bar{Z}_{i}=Y_{ki}$, then it holds that
$\mathbb{E}[v_{i}|T_{i}]$ $=\sum_{a=0}^{1}1\{Y_{ki}=a\}\mathbb{E}\left[\cdot\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}|T_{i}\right],$
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $11\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=$ 1, then it holds that
@@ 5 Theory
In this section, we present consistency and coverage results for the general class of Hájek-normalized estimators described in Section 4. Specifically, we show that our point estimate $\hat{\tau}$ is consistent for a lower bound on the estimand while equaling a Hájek-weighted contrast; we show that the variance estimators $V_{k}$ concentrates and is either unbiased or conservative for the variance that is required for our interval estimates to cover; and we show that our intervals have the desired asymptotic coverage, albeit with a technical modification to cover degenerate settings in which normality does not hold.
To describe our requirements on the experiment design, for $i\in[N]$ let $\Gamma_{i}\subseteq[N]$ denote the smallest subset of units whose treatment assignment is sufficient to determine the values of $T_{i},Z_{i}$, and $1\{X\in R_{i}\}$ for all $X\in\{0,1\}^{N}$. We will call $\Gamma_{i}$ the design neighborhood of unit $i$. We stress that it places no restrictions on the outcome model or interference.
Assumption 1 describes the experiment design that we study. It assumes a sequence of experiments in which the design neighborhoods are bounded in size and overlap, Bernoulli randomization with treatment probabilities bounded away from 0 and 1, and places a positivity condition on $P(\bar{Z}_{i}|T_{i})$ under the hypothesis that $\phi_{i}^{*}=1$. While many experiment designs are possible, we study independent Bernoulli randomization as a starting point, leading to the simplest proofs of consistency and asymptotic normality.
@@
The treatment assignments $\{X_{i}\}_{i=1}^{N}$ are indpendent Bernoulli random variables, which satisfy for fixed constants $d_{\max}$ and $p_{\min}>0$ as $N\rightarrow\infty$ that
$|\Gamma_{i}|$ $\leq d_{\max}$ $\forall\ i\in[N]$ (40)
$\sum_{j=1}^{N}1\{x_{i}\in\Gamma_{j}\}$ $\leq d_{\max}$ $\forall\ i\in[N]$ (41)
$p_{\min}\leq P(X_{i}=1)$ $\leq 1-p_{\min}$ $\forall\ i\in[N]$ (42)
and for all $i:\phi_{i}^{*}=1$, it holds for $\bar{Z}$ given by (39) that
[…]
The support condition (43) requires $T_{i}$, $Z_{i}$, and $R_{i}$ to be chosen so that the entries of $X$ needed to determine $Z_{i}$ and membership in $R_{i}$ do not fully overlap with those needed to determine $T_{i}$. For example, if $T_{i}=X_{\eta_{i}}$, then $Z_{i}$ and $R_{i}$ should depend at least partially on units outside of $\eta_{i}$.
Theorem 4 is a consistency result. It states that $\hat{\Phi}^{\textup{Haj}}_{k}$ converges to $\Phi$, and that $\hat{\tau}$ is asymptotically a lower bound for $\tau$, and equals a Hájek weighted contrast between units, given by (46). Table 2 gives examples of (46) for the estimands considered in the paper.
@@
Let Assumption 1 hold. Then
$\hat{\Phi}_{k}^{\textup{Haj}}=\Phi+O_{P}(N^{1/2}),\qquad k=1,2$ (44)
and for $\hat{\tau}=N-\min\left\{\max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\ ,\ \max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right\}$ it holds that
$\hat{\tau}\leq\tau+O_{P}(N^{1/2}).$ (45)
For arbitrary experiment designs (i.e., no longer requiring Assumption 1), it holds that
$\hat{\tau}=\left|\frac{1\{X\in R_{i},Z_{i}=1\}N}{P(X\in R_{i},Z_{i}=1|T_{i})\hat{N}_{1}}Y_{i}-\sum_{i=1}^{N}\frac{1\{X\in R_{i},Z_{i}=0\}N}{P(X\in R_{i},Z_{i}=0|T_{i})\hat{N}_{0}}Y_{i}\right|,$ (46)
where $\hat{N}_{a}=\displaystyle{\sum_{i=1}^{N}}\dfrac{1\{X\in R_{i},Z_{i}=a\}}{P(X\in R_{i},Z_{i}=a|T_{i})}$.
Theorem 5 pertains to variance estimation. It states that $V_{k}$ converges to an upper bound on the variance of $\hat{\Phi}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
@@
Let Assumption 1 hold. Then for any $\delta>0$ it holds that
$V_{k}=\mathbb{E}[V_{k}]+O_{P}(N^{1/2+\delta}),\qquad k=1,2.$ (47)
@@
$\mathbb{E}V_{k}\geq{\operatorname{Var}\,}\hat{\Phi}_{k}^{\textup{Haj}},$ (48)
with equality if the exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant and
$P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $a,b\in\{0,1\}$ and $i,j\in[N]^{2}$.
Theorem 6 gives coverage properties for the proposed confidence interval given by (30). It states that if the variance estimators $V_{k}$ are lower bounded by a vanishing fraction of $N$, then the intervals will have at least the desired asymptotic coverage level.
@@
Let Assumption 1 hold. Let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{V_{k}(\phi)\ ,\ \frac{1.01\ N^{2/3+0.01}}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\right\}$
and let $CI$ denote
$CI=N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (49)
Then $\tau\geq CI$ with probability converging to at least $1-\alpha$.
All theorems are proven in the appendix.
@@ 6.1 Simulation Study
where the parameters $(u_{i},\alpha_{0i},\alpha_{1i},\alpha_{2i})$ were randomly generated, with $u_{i}\sim\textup{Unif }[0,1]$, and $\alpha_{0i},\alpha_{1i},\alpha_{2i}$ nonnegative. For simulations involving $\tau^{\textup{nonneighbors}}$, the outcome model was changed to
$Y_{i}=1\{\textup{if }u_{i}\leq\alpha_{0i}+\alpha_{1i}\cdot X_{i}+\alpha_{2i}\cdot\bar{W}_{i}\}$ (51)
@@
where $\bar{W}_{i}$ denotes the thresholded number of treated units in $\eta_{i}\cup\eta_{i}^{(2)}$ excluding unit $i$, so that neighbors in $\eta_{i}$ and $\eta_{i}^{(2)}$ were equally important in determining the outcome of unit $i$. Unit $i$ was considered to be affected by treatment if $Y_{i}$ varied as a function of $X$ under the generated parameter values, and was considered to be affected by treatment of others if $Y_{i}$ varied as a function of $X_{-i}$ for any value of $X_{i}$, and so forth.
@@
Hájek normalization was used on all estimators. Estimation of $\tau^{\textup{basic}}$ involved both $W$ and $X$, following (14) with Hájek normalization. For comparison of coverage rates and interval widths, the treatment effect
$EATE=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|X\in R_{i},Z_{i}=1]-\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|X\in R_{i},Z_{i}=0],$
@@
was also estimated, for $\{R_{i}\}$ and $\{Z_{i}\}$ given by Table 1 for each estimand. For example, when estimating $\tau^{\textup{indirect}}$, we compared to the indirect effect given by
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|W_{i}=1]-\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[Y_{i}|W_{i}=0].$
@@
For estimation and inference of these treatment effects, Hájek normalization and conservative variance estimates were used following Gao and Ding, (2023).
@@
The results of the simulations are the following. For the estimands $\tau^{\textup{basic}}$, $\tau^{\textup{tr}}$, and $\tau^{\textup{control}}$, point estimates were unbiased (as might be expected given that the unit effects were monotone functions of the contrast variables), and the confidence intervals exhibited approximate calibration of coverage. For these estimands, the exposure mapping $T_{i}$ is constant, allowing for unbiased variance estimation as discussed in Theorem 5. For the remaining estimands, estimation was conservatively biased, and interval coverage rates exceeded the desired level. Table 3 also shows that the interval widths for our estimands were competitive, and often smaller than, those of comparable EATE treatment effects. Coverage of the EATE interval estimates was conservative, always exceeding 99%.
Actual Bias of RMSE of Coverage Avg CI Avg Width
Estimand N Value$/N$ Estimator Estimator of 95% CI Width of EATE CI
$\tau^{\textup{basic}}$ 250 0.40 -0.00 0.09 0.95 0.18 0.27
500 0.36 -0.00 0.06 0.95 0.12 0.22
$\tau^{\textup{tr}}$ 250 0.21 -0.00 0.11 0.94 0.20 0.29
500 0.16 -0.00 0.06 0.95 0.12 0.23
$\tau^{\textup{control}}$ 250 0.21 -0.00 0.10 0.93 0.20 0.28
500 0.18 -0.00 0.06 0.94 0.12 0.23
$\tau^{\textup{indirect}}$ 250 0.36 0.14 0.16 0.99 0.17 0.21
500 0.30 0.12 0.13 1.0 0.10 0.16
$\tau^{\textup{nonneighbor}}$ 250 0.36 0.15 0.18 1.0 0.27 0.21
500 0.35 0.14 0.15 1.0 0.24 0.17
@@
Table 3: Results of Simulation Study. Bold values indicate approximate 95% coverage, or that CI widths (measured as distance from point estimate to interval boundary) were smaller than those of comparable treatment effect.

### v4

@@ 4.3 Variance Estimation and Inference for General Case
$v_{i}=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}$ (34)
so that $\hat{\tau}_{k}^{\textup{Haj}}=N-\sum_{i\in{\mathcal{I}}}v_{i}$. It follows that the variance of $\hat{\tau}_{k}^{\textup{Haj}}$ can be written as a sum of covariances
${\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}(v_{i},v_{j})$
Unlike the special case where $T_{i}$ is constant for all units, generally we cannot compute ${\operatorname{Cov}}(v_{i},v_{j})$ exactly without knowledge of the outcome models $f_{i}$ and $f_{j}$, even if $i,j\in{\mathcal{I}}$. As a result, the covariance terms must be estimated. We thus propose variance estimator $\hat{V}_{k}$ given by
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}\Bigg\{\mathbb{E}\left[v_{i}v_{j}|T_{i},T_{j}\right]-\mathbb{E}\left[v_{i}|T_{i}\right]\cdot\mathbb{E}\left[v_{j}|T_{j}\right]\cdot\\ \frac{P(T_{i})\cdot P(T_{j})}{P(T_{i},T_{j})}\Bigg\}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})}$ (35)
@@
$\hat{V}_{k}=\sum_{i\in{\mathcal{I}}}\sum_{j\in{\mathcal{I}}}{\operatorname{Cov}}(v_{i},v_{j})\cdot\frac{1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}}{P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj})},$
@@
If $\hat{\tau}_{k}^{\textup{Haj}}$ for $k=1,2$ are consistent and asymptotically normal, with variance consistently (or conservatively) estimated by $\hat{V}_{k}$, then we can construct an asymptotic $(1-\alpha)$ level one-sided confidence interval for $\tau$, with lower boundary given by
$\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{2}(\phi)}\,\right\},$ (36)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ denote $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$, evaluated under the hypothesis that ${\mathcal{I}}=\{i:\phi_{i}=1\}$.
To evaluate the terms $\hat{\tau}_{k}^{\textup{Haj}}(\phi)$ and $\hat{V}_{k}(\phi)$ appearing in (36), we may use Propositions 4 and 5. Proposition 4 shows that the sampling weights $P(\bar{Z}_{i}=Y_{ki}|T_{i})$ appearing in $\hat{\tau}_{k}^{\textup{Haj}}$ given by (31) do not require knowledge of the outcome mappings $\{f_{i}\}_{i=1}^{N}$. Proposition 5 gives computable expressions for the terms appearing in $\hat{V}_{k}$ given by (35).
Proposition 4.
If $i\in{\mathcal{I}}$ and $X\in R_{i}$, then it holds that
$P(\bar{Z}_{i}=Y_{ki}|T_{i})=Y_{ki}\cdot P(\bar{Z}_{i}=1|T_{i})+(1-Y_{ki})\cdot P(\bar{Z}_{i}=0|T_{i}).$
Proposition 5.
If $i\in{\mathcal{I}}$ and $\bar{Z}_{i}=Y_{ki}$, then it holds that
$\mathbb{E}[v_{i}|T_{i}]$ $=\sum_{a=0}^{1}1\{Y_{ki}=a\}\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}N}{P(\bar{Z}_{i}=a|T_{i})\hat{N}_{a}}|T_{i}\right],$
If $i,j\in{\mathcal{I}}$ and $1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=1$, then it holds that
@@ 5 Theory
In this section, we give consistency and coverage results for the general class of Hájek-normalized estimators described in Section 4. Specifically, we show that $|\Delta^{\textup{Haj}}|$ asymptotically lower bounds the estimand; we show that the variance estimator $\hat{V}_{k}$ concentrates and is either unbiased or conservative; and we show that our intervals have the desired asymptotic coverage, albeit with a technical modification to cover settings approaching degeneracy, in which case normality may not hold.
To describe our requirements on the experiment design, for $i\in[N]$ let $\Gamma_{i}\subseteq[N]$ denote a subset of units whose treatment assignment is sufficient to determine the values of $T_{i}$ and $\bar{Z}_{i}$ for all $X\in\{0,1\}^{N}$, so that for all $X,X^{\prime}\in\{0,1\}^{N}$ satisfying $X_{\Gamma_{i}}=X_{\Gamma_{i}}^{\prime}$, it holds that
$T_{i}(X)=T_{i}(X^{\prime}),\textup{ and }\bar{Z}_{i}(X)=\bar{Z}_{i}(X^{\prime})$
We will call $\Gamma_{i}$ the design neighborhood of unit $i$. We stress that requirements on $\{\Gamma_{i}\}_{i=1}^{N}$ restrict the experiment designs and estimands which may be explored, but places no assumptions on the outcome models $\{f_{i}\}_{i=1}^{N}$ that describe interference, which are unknown and uncontrolled by the researcher.
Assumption 1 describes a simple experiment design that we study. It assumes a sequence of experiments in which the design neighborhoods are bounded in size and overlap, Bernoulli randomization with treatment probabilities bounded away from 0 and 1, and places a positivity condition on $P(\bar{Z}_{i}|T_{i})$.
@@
The treatment assignments $\{X_{i}\}_{i=1}^{N}$ are independent Bernoulli random variables, which satisfy for fixed constants $p>0$ and $d>0$ as $N\rightarrow\infty$ that
$|\Gamma_{i}|$ $\leq d$ $\forall\ i\in[N]$ (37)
$\sum_{j=1}^{N}1\{i\in\Gamma_{j}\}$ $\leq d$ $\forall\ i\in[N]$ (38)
$p\leq P(X_{i}=1)$ $\leq 1-p$ $\forall\ i\in[N]$ (39)
$P(\bar{Z}_{i}=a\,|\,T_{i}=t)$ $\geq p$ $\forall\ i\in[N],a\in\{0,1\},t\in\mathcal{T}_{i}$ (40)
where $\mathcal{T}_{i}$ denotes the support of $T_{i}$.
The positivity condition (40) requires $T_{i}$, $Z_{i}$, and $R_{i}$ to be chosen so that $Z_{i}$ and $R_{i}$ are not deterministic conditional on $T_{i}$. For example, if $T_{i}=X_{\mathcal{N}_{i}}$, then $Z_{i}$ and $R_{i}$ should depend at least partially on units outside of $\mathcal{N}_{i}$.
Theorem 1 is a consistency result. It states that $\hat{\tau}^{\textup{Haj}}_{k}$ converges to $\tau$, and that the absolute value of $\Delta^{\textup{Haj}}$, the Hájek weighted contrast between units given by (33), is asymptotically a lower bound for $\tau$.
@@
It holds that
$|\Delta^{\textup{Haj}}|$ $\leq\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ (41)
If Assumption 1 holds, then
$\hat{\tau}_{k}^{\textup{Haj}}=\tau+O_{P}\left(N^{1/2}\right),\qquad k=1,2$ (42)
in which case (41) implies $|\Delta^{\textup{Haj}}|\leq\tau+O_{P}\left(N^{1/2}\right)$
Theorem 2 pertains to variance estimation. It states that $\hat{V}_{k}$ concentrates to its expectation which upper bounds the variance of $\hat{\tau}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
@@
Let Assumption 1 hold. It holds that
$\hat{V}_{k}=\mathbb{E}[\hat{V}_{k}]+O_{P}\left(N^{1/2}\log N\right),\qquad k=1,2.$ (43)
@@
$\mathbb{E}\hat{V}_{k}\geq{\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}},$ (44)
with equality if all of the following conditions hold:
  1. 1.
    The exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant
  2. 2.
    $P(\bar{Z}_{i}=a)>0$ for all $i\in[N]$ and $a\in\{0,1\}$
  3. 3.
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $i,j\in[N]$ such that $i\neq j$, and $a,b\in\{0,1\}$.
Theorem 3 gives coverage properties for the proposed confidence interval given by (36). It states that if the variance estimators $\hat{V}_{k}$ are lower bounded by a vanishing fraction of $N$, then the interval will have at least the desired asymptotic coverage level.
@@
Let Assumption 1 hold. For any $\epsilon>0$, let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{\hat{V}_{k}(\phi)\ ,\ \frac{N^{2/3+\epsilon}}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\right\}$ (45)
and let $LB_{1-\alpha}$ denote
$LB_{1-\alpha}=\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (46)
Then $\tau\geq LB_{1-\alpha}$ with probability converging to at least $1-\alpha$.
In practice, we expect usage of $\tilde{V}_{k}(\phi)$ rather than $\hat{V}_{k}(\phi)$ to make little or no difference, as the two terms are equal except in cases involving interval half-widths of order $O(N^{1/3})$ or smaller. This is a vanishing fraction of the usual $O(\sqrt{N})$ rate that arises for estimation of population sizes under simple random sampling, and we do not expect the interval widths found by (46), which are widened by the optimization over $\phi$, to be so small in practice.
Additional Experiment Designs
We have considered independent Bernoulli randomization primarily as a starting point for its simplicity. In the supplemental materials, we generalize the theorems to include two-stage experiment designs, in which the size of the design neighborhoods may grow with $N$. The result will allow contrasts (i.e., $\bar{Z}$) involving not only the treatment of each unit and a bounded number of close friends, but also the number of treated units in some broader division such as geographic region, or overlapping collections of neighboring geographic regions, in which the region sizes may grow at rate $O(N^{1-\delta})$ for any $\delta>0$.
@@ 6.1 Simulation Study
where the scalar parameters $u_{i},\alpha_{0i},\alpha_{1i},\alpha_{2i}$ were randomly generated, with $u_{i}\sim\textup{Unif }[0,1]$, and $\alpha_{0i},\alpha_{1i},\alpha_{2i}$ nonnegative. For simulations involving $\tau^{\textup{nonneighbors}}$, in (47) we change $W_{i}$ to $\bar{W}_{i}$, the thresholded number of treated units in ${\mathcal{N}}_{i}\cup{\mathcal{N}}_{i}^{(2)}$ excluding unit $i$, so that units in ${\mathcal{N}}_{i}$ and ${\mathcal{N}}_{i}^{(2)}$ were equally important in determining the outcome of unit $i$. To compute the true value of the estimand, unit $i$ was considered to be affected by treatment if $Y_{i}$ varied as a function of $X$ under the generated parameter values, and was considered to be affected by treatment of others if $Y_{i}$ varied as a function of $X_{-i}$ for any value of $X_{i}$, and so forth.
@@
Estimation of $\tau^{\textup{basic}}$ used the contrast variable $Z$ given by (9). Hájek normalization was used on all estimators. For comparison, the treatment effect
$EATE=\frac{1}{N}\sum_{i=1}^{N}\Big(\mathbb{E}[Y_{i}|\bar{Z}_{i}=1]-\mathbb{E}[Y_{i}|\bar{Z}_{i}=0]\Big),$
@@
was also estimated, where $\bar{Z}$ followed Table 1 for each estimand. For example, when estimating $\tau^{\textup{indirect}}$, we compared to the indirect effect given by
$\frac{1}{N}\sum_{i=1}^{N}\Big(\mathbb{E}[Y_{i}|W_{i}=1]-\mathbb{E}[Y_{i}|W_{i}=0]\Big).$
@@
For estimation and inference of these treatment effects, Hájek normalization and conservative variance estimates were used following Gao and Ding, (2023). [^5] Variance estimates for EATE followed Eq. (11) (“adjusted HAC covariance estimator”) of Gao and Ding, (2023) (arxiv version v3), using our $\bar{Z}_{i}$ for their treatment condition $T_{i}$. For the simulations, the correctly specified bandwidth $b_{n}=2$ was used.
@@
The results of the simulations are the following. For the estimands $\tau^{\textup{basic}}$, $\tau^{\textup{tr}}$, and $\tau^{\textup{control}}$, point estimates were unbiased (as might be expected given that the unit effects were nonnegative, consistent with the intuition of Example 1), and the confidence intervals exhibited approximate calibration of coverage. For these estimands, the exposure mapping $T_{i}$ is constant, allowing for unbiased variance estimation as discussed in Theorem 2. For the remaining estimands, estimation was conservatively biased, and interval coverage rates were 100%. Table 3 also shows that the interval half-widths for our estimands were competitive, and often smaller than, those of comparable EATE treatment effects. Coverage of the EATE interval estimates was conservative, always exceeding 99%.
Our intuition for the smaller intervals of our approach is the following. Whereas treatment effect estimation uses $\Delta^{\textup{Haj}}$ as a point estimate and therefore must control its variability, we use $\Delta^{\textup{Haj}}$ rather as a lower bound for the point estimators $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$. Thus, to control anticonservativeness we need to consider the variability of these point estimators, which we expect to be less than that of $\Delta^{\textup{Haj}}$ and also easier to estimate, as only units in $\mathcal{I}$} are considered.
Actual $|\Delta^{\textup{Haj}}|/N$ Coverage CI Half- $\widehat{EATE}$ EATE CI
Estimand N Value/N/N Avg SD of 95% CI width Avg SD Half-width
$\tau^{\textup{basic}}$ 250 0.43 0.43 0.08 0.94 0.16 0.43 0.08 0.26
500 0.37 0.37 0.05 0.96 0.11 0.36 0.05 0.21
$\tau^{\textup{tr}}$ 250 0.22 0.22 0.08 0.95 0.15 0.22 0.08 0.26
500 0.18 0.18 0.05 0.95 0.11 0.18 0.05 0.20
$\tau^{\textup{control}}$ 250 0.23 0.23 0.10 0.95 0.16 0.23 0.10 0.29
500 0.21 0.21 0.06 0.96 0.12 0.21 0.06 0.23
$\tau^{\textup{indirect}}$ 250 0.37 0.22 0.06 1.0000 0.15 0.22 0.06 0.20
500 0.33 0.20 0.04 1.0000 0.10 0.19 0.04 0.16
$\tau^{\textup{nonneighbor}}$ 250 0.37 0.22 0.07 1.0000 0.21 0.22 0.06 0.20
500 0.33 0.20 0.04 1.0000 0.18 0.19 0.04 0.16
@@
Table 3: Results of Simulation Study. Bold values indicate approximate 95% coverage, or that CI half-widths were smaller than those of comparable treatment effect. CI Half-width defined as average distance from point estimate to the interval boundary. For our estimands, $|\Delta^{\textup{Haj}}|/N$ was point estimate. 500 trials per simulation.